OpenStudy (anonymous):

when a polynomial is divided by x-4, the quotient is x^2-5x+4+1/x+4. what is the polynomial?

OpenStudy (anonymous):

i think multiply (x-4) by (x^2-5x+4+1) and just add on the x+4 so (x-4)*(x^2-5x+4+1) + x+4

OpenStudy (anonymous):

could be wrong

OpenStudy (cruffo):

Let y = unknown polynomial Then y/(x+4) = x^2-5x+4 + 1/(x+4) Multiply both sides by (x+4) to solve for y

OpenStudy (anonymous):

i suck at this.

OpenStudy (cruffo):

Or do something close to what @timo86m did (small adjustment) (x-4)*(x^2-5x+4) + 1

OpenStudy (cruffo):

This follows the pattern polynomial = divisior * quotient + remainder

OpenStudy (anonymous):

so would the answer be 20x^4-15?

OpenStudy (cruffo):

That's not what I got. Let me check my work.

OpenStudy (cruffo):

Question: In the problem: \[ \large x^2-5x+4+\frac{1}{x+4}\] Is the remainder over x+4 or x-4. If your dividing by x-4, the denominator should be x-4 as well. Would you check that for me?

OpenStudy (anonymous):

it's over x-4 sorry

OpenStudy (cruffo):

Cool.

OpenStudy (anonymous):

so my answer wasn't right ?

OpenStudy (cruffo):

I got something different.

OpenStudy (anonymous):

so the remainder is -15?

OpenStudy (cruffo):

The "remainder" is the 1 that was originally divided by x-4.

OpenStudy (anonymous):

ok so i don't have to put that then i just have to x^3-9x^2+24x-15

OpenStudy (cruffo):

Yep. And it checks out. Just divide the polynomial you found by x-4.

OpenStudy (anonymous):

so my ploynomiial is x^3-9x^2+24-15 or x^2-5x+4??

OpenStudy (cruffo):

polynomial = x^3-9x^2+24x-15 divisor = x - 4 quotient = x^2-5x+4 remainder = 1

OpenStudy (anonymous):

can i ask u another one?

OpenStudy (cruffo):

Sure, post a new question. I'll look for you.